OpenStudy (anonymous):

You are going to fence an area that is in the shape of a right triangle. The two legs measure 7.8 m and 5.78 m. Estimate the meters of fencing you will need for the longest side, using whole meters and square root.

OpenStudy (smileyxl3):

Since its a right triangle a^2 + b^2 = c^2

OpenStudy (anonymous):

Are you familiar with the Pythagorean theorem?

OpenStudy (anonymous):

^

OpenStudy (smileyxl3):

Lol great minds think alike

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[a^2+b^2=c^2 \] where a and b are shorter sides of the triangle, and c is the longest side.

OpenStudy (smileyxl3):

You have to round them up too. 7.8=8 and 5.78=6. that should be easier

OpenStudy (anonymous):

So for this we can represent this as \[(5.78)^2+(7.8)^2=c^2\]

OpenStudy (anonymous):

So\[c=\sqrt{(5.78)^2+(7.8)^2}\]

OpenStudy (anonymous):

right

OpenStudy (smileyxl3):

\[6^2+8^2=c^2\]

OpenStudy (anonymous):

\[c \approx100m\]

OpenStudy (smileyxl3):

\[\sqrt{36+64}\]

OpenStudy (anonymous):

Oops...

OpenStudy (anonymous):

Forgot my square root!!

OpenStudy (smileyxl3):

@escolas I think it would have been easier to calculate if you just rounded the sides first. and yea you forgot your square

OpenStudy (anonymous):

\[c \approx \sqrt{100}\]

OpenStudy (smileyxl3):

perimeter = 10 + 8 + 6 = 24. By the way, this is a 3 4 5 triangle ;)

OpenStudy (anonymous):

It would have been, but for real math it's much less accurate. I mean I wasn't sure it'd come out to 9 or 10, it did happen to be barely closer to 10 so it would round up.